OpenStudy (anonymous):

33 bigmommatiad 0we are doing a different problen hr it is the number of state and federal inmates in millions during year x, were x>2002 can be approx by following formula n=0.05x-98.65 determine year in which there were 1.5 million inmates 6 minutes ago

OpenStudy (anonymous):

nubeer I opened a new one my system was being funny please still help me

OpenStudy (anonymous):

(n-98.65)/.05x so would that be omg I do not know

OpenStudy (nubeer):

hmm u dont have to put x with 0.05 as u have made x the subject and u wrote it wrong its (n+98.65)/.05=x where n is 1.5

OpenStudy (anonymous):

so the answer would be 2003 then right

OpenStudy (nubeer):

yes.

OpenStudy (anonymous):

ok I did this question but I want to check my work 4.3 x 10^5 -------------- 6.2 x 10^7 divide 4.3/6.2 I get .69 then .69 x 10^2 then I get 69 it is wanting the total family networth dollars in this would it be 690,000

OpenStudy (anonymous):

for 1 year

OpenStudy (nubeer):

\[\frac{ 4.3 * 10^{5} }{ 6.2 *10^{7} }\] \[\frac{ 10^{5} }{ 10^{7}}\] \[\frac{ 1 }{ 10^{2} }\]

OpenStudy (nubeer):

can you see your mistake?

OpenStudy (anonymous):

how did you get the 1/10^2, I got the 10^2 but from what I read it is saying to take the decinmal on the .69 over 2

OpenStudy (nubeer):

then can u tell me how you got 10^2?

OpenStudy (anonymous):

10^7-10^5

OpenStudy (nubeer):

ok and is 10^7 is in denominator or numerator?

OpenStudy (anonymous):

anyone else want to help I appreicate it I have been doing this math now for 9 hours and still have a lot of problem to go

OpenStudy (anonymous):

numerator I believe

OpenStudy (nubeer):

hmm well see the question again 10^7 is in denominator.

OpenStudy (anonymous):

How do you get this math i am going crazy trying to figure it out

OpenStudy (nubeer):

lol nothing much just use of brain :P

OpenStudy (anonymous):

oh so when it is at bottom you will put a 1/10^2

OpenStudy (nubeer):

yes.. and then?

OpenStudy (anonymous):

so did I do the first portaion right with 4.3/6/2

OpenStudy (nubeer):

yes that was right.

OpenStudy (anonymous):

ok so if you have the .69 you would have 69/1 then?

OpenStudy (anonymous):

or 1/69

OpenStudy (nubeer):

hmmm why dont u bring 1/10^2 above means liek this 0.69 * 10^-2 )or u can write (69/10) * (1/10^2)

OpenStudy (anonymous):

I want to thank you for helping me, I did all of the homewerk for these and did well but when it came to the test for the first 8 weeks I was lost, I printed all of the questions on the test and also printed all of the chapters and I am still lost

OpenStudy (anonymous):

you lost me on the 1/10^2

OpenStudy (nubeer):

well it's not hard.. u just need more practise.. u should solve all these questions agian yourself without any help..

OpenStudy (anonymous):

so to get my answer I would take the e1/10*2 then times that with the 6.9

OpenStudy (nubeer):

yes and it's 0.69 not 6.9 :P if i am nto wrong :p

OpenStudy (anonymous):

yes you are correct, so that would be 1/10*2=.1*2*.69 right

OpenStudy (anonymous):

so I came out with .0069

OpenStudy (nubeer):

yes that should be